# Embedding Self-Correction as an Inherent Ability in Large Language Models for Enhanced Mathematical Reasoning

### v1

@@ Abstract
Accurate mathematical reasoning with Large Language Models (LLMs) is crucial in revolutionizing domains that heavily rely on such reasoning. However, LLMs often encounter difficulties in certain aspects of mathematical reasoning, leading to flawed reasoning and erroneous results. To mitigate these issues, we introduce a novel mechanism, the Chain of Self-Correction (CoSC), specifically designed to embed self-correction as an inherent ability in LLMs, enabling them to validate and rectify their own results. The CoSC mechanism operates through a sequence of self-correction stages. In each stage, the LLMs generate a program to address a given problem, execute this program using program-based tools to obtain an output, subsequently verify this output. Based on the verification, the LLMs either proceed to the next correction stage or finalize the answer. This iterative self-correction process allows the LLMs to refine their reasoning steps and improve the accuracy of their mathematical reasoning. To enable the CoSC mechanism at a low cost, we employ a two-phase finetuning approach. In the first phase, the LLMs are trained with a relatively small volume of seeding data generated from GPT-4, establishing an initial CoSC capability. In the second phase, the CoSC capability is further enhanced by training with a larger volume of self-generated data using the trained model in the first phase, without relying on the paid GPT-4. Our comprehensive experiments demonstrate that CoSC significantly improves performance on traditional mathematical datasets among existing open-source LLMs. Notably, our CoSC-Code-34B model achieved a 53.5% score on MATH, the most challenging mathematical reasoning dataset in the public domain, surpassing the performance of well-established models such as ChatGPT, GPT-4, and even multi-modal LLMs like GPT-4V, Gemini-1.0 Pro, and Gemini-1.0 Ultra. It’s important to note that, unlike these proprietary models, our CoSC performs inference in a zero-shot manner, without the need for demonstrations. To the best of our knowledge, this is the first work in the literature to embed self-correction as an inherent capability in large language models and achieve significant performance improvements. To facilitate open-source research on improving the reasoning capabilities of LLMs, the code and data for this work will be released upon the acceptance of this paper.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 (OpenAI, 2023), have recently demonstrated state-of-the-art performance across a variety of natural language processing (NLP) tasks, including natural language generation and understanding (Chowdhery et al., 2023; Team et al., 2023; Anil et al., 2023; Penedo et al., 2023). Despite their success, LLMs often struggle with mathematical reasoning tasks due to their lack of explicit logical reasoning and judgment, which are crucial for solving such problems. Moreover, there is a fundamental gap between natural language and the language of mathematical formulas, which further complicates these tasks. As a result, accurate mathematical reasoning remains an essential yet challenging capability for LLMs to develop, in order to further advance various domains. Consequently, it is still an open challenge to tackle mathematical problems for existing open-source LLMs, such as LLaMA-2 (Touvron et al., 2023).
@@
Most current methods (Yu et al., 2023; Gou et al., 2023b) generate finetuning datasets by prompting GPT-4 to rephrase mathematical questions from various perspectives or incorporate chain-of-thoughts (CoT) analysis (Wei et al., 2022) and program-of-thoughts (PoT) code (Chen et al., 2023a) to diversify mathematical responses. As a result, LLMs trained on these datasets can comprehend different questions and learn to use code to solve mathematical problems. However, precise multi-round reasoning capabilities remain challenging for them. Once potential errors occur throughout reasoning stages, it can lead to incorrect results. Consequently, it is essential to incorporate a self-correction mechanism into mathematical responses, which can enable LLMs to learn to correct themselves in multiple rounds. In addition, while supervised finetuning datasets can be developed with GPT-4’s assistance, it still requires human experiments with different prompts and the cost of using the interface. This highlights the need for effective finetuning with unlabeled datasets.
Figure 1: Comparison of four reasoning frameworks for solving an example mathematical question. (a) Chain of Thoughts (CoT) (Wei et al., 2022). (b) Program of Thoughts (PoT) (Chen et al., 2023a). (c) ToRA (Gou et al., 2023b) that incorporates CoT, PoT, and the utilization of tools. (d) Our proposed CoSC consisits of a sequence of multiple self-correction stages (two stages are shown in this example). Each stage has four sub-stages: (p1) LLMs generate program w.r.t. the question; (o1) execute the program to obtain program output; (v1) perform two-step verification for consistency of the question with both the generated program and the program output; (c1) conclude a refined answer or continue the next subsequent self-correction stage depending on the verification result. The final answer is extracted from the last conclusion sub-stage with regular expression matching.
@@
Utilizing GPT-4 can be expensive, especially when dealing with large volumes of training data. Alternatively, we propose a cost-free method to further boost performance through self-enhancement in the second phase. This second phase, referred to as CoSC self-enhancement, builds upon the foundational learning by further adapting the LLMs obtained from the first phase with self-generated trajectories. These trajectories are produced by the models trained in the foundational phase, allowing for the generation of a substantial volume of data without additional GPT-4 intervention. In both phases, we retain only the trajectories whose answers match the ground-truth labels of the corresponding questions.
@@
  • To implement the CoSC mechanism at a low cost, we introduce a two-phase finetuning approach. The first phase involves CoSC foundational learning, where we use a relatively small volume of seeding data generated by GPT-4. In the second phase, CoSC self-enhancement occurs using a larger volume of self-generated data with the model obtained from the first phase, without relying on paid GPT-4.
  • Our comprehensive experimental evaluations demonstrate that the CoSC mechanism provides a new benchmark for performance on established mathematical datasets when compared to existing open-source LLMs. Notably, our CoSC-Code-34B model achieves superior performance over both closed-source LLMs and some of multi-modal LLMs, particularly on the challenging MATH dataset. These findings underscore the potential of our CoSC to revolutionize the accuracy and reliability of LLMs in complex problem-solving domains.
  • The proposed CoSC mechanism, by embedding self-correction as an inherent capability in LLMs, enables them to think before responding to a question, creating an internal chain of self-correction to progressively verify and rectify their original answers. It is more akin to the slow thinking process of humans, which is particularly helpful in solving difficult mathematical reasoning problems. This approach can provide valuable insights for future research and contribute to the ongoing advancement of LLMs across various domains.
@@ 2.1 LLMs for Mathematical Reasoning
Mathematical reasoning (Liu et al., 2023a; Wang et al., 2023a; Huang et al., 2024b; Toshniwal et al., 2024; Chen et al., 2024a; Zhang et al., 2024) is a challenging reasoning task for LLMs, which requires the ability to understand mathematical concepts, computation and multi-round reasoning. Existing mathematical reasoning approaches can be broadly classified into three categories: (1) Prompting methods (Wei et al., 2022; Chen et al., 2023a; Wang et al., 2023b) focus on extracting the inherent mathematical reasoning skills of LLMs by utilizing well-crafted prompting strategies during inference. Notably, they leverage the existing knowledge in LLMs without the need for parameter updates. (2) Pretraining methods (Azerbayev et al., 2023; Fu et al., 2023) pre-train LLMs on large-scale corpora containing mathematical problems and related content with language modeling objectives. The goal is to train a base foundation language model as a platform for mathematical domain. (3) Finetuning methods (Luo et al., 2023; Yu et al., 2023; Yue et al., 2023; Gou et al., 2023b; Liu et al., 2023b; Gao et al., 2024) refine the mathematical reasoning capabilities of LLMs by offering more targeted training. Central to this approach is the generation of high-quality question-response pair datasets, which needs the assistance of complexity-based chain-of-thoughts prompting or tools-based augmentation. In this paper, we propose a Chain of Self-Correction (CoSC) along this direction that incorporates an iterative self-correction process into datasets generation.
@@ 2.2 Existing methods related to Self-Correction in LLMs
Mathematical reasoning poses a significant challenge due to its demand for precise multi-round logical reasoning to solve problems. The potential for errors increases with each reasoning step, making it crucial for LLMs to have the ability to self-correct in order to produce accurate results. There are some recent studies (Chen et al., 2023b; Gou et al., 2023a; Lightman et al., 2023; Huang et al., 2024a; Chen et al., 2024b) attempt to enable large language models to perform self-correction by either prompting methods or or fine-tuning methods. For prompting methods, they can correct their responses when interacting with external tools (Gou et al., 2023a; Chen et al., 2023b), such as search engines and calculators, or designing complex prompts (Chen et al., 2024b). Notably, prompting methods can be orthogonally combined with finetuning methods. For finetuning methods, previous works (Yu et al., 2024; An et al., 2023) only model single-round correction during the training stage, and perform verification in a straightforward manner during the inference stage.
@@ 2.3 Difference between the proposed method and existing self-correction techniques
The existing self-correction research can be broadly classified into two categories: prompt-based methods (Gou et al., 2023a; Chen et al., 2023b; 2024b) and SFT-based methods (Yu et al., 2024; An et al., 2023). Our approach falls under the SFT-based category and fundamentally differs from prompt-based methods. Specifically, prompt-based methods leverage the intrinsic capabilities of large language models (LLMs) for mathematical reasoning. In contrast, our method embeds the chain of self-correction, a strong reasoning mechanism, as an inherent capability in the LLMs through parameter fine-tuning. Furthermore, compared to existing SFT-based methods, our approach models mathematical reasoning as a multi-round procedure during the training stage. In addition, these SFT-based works (Yu et al., 2024; An et al., 2023) perform the verification process in a straightforward manner, whereas our work conducts verification in a step-by-step manner by generating intermediate verification steps. Similar to the essence of the CoT approach, our model first verifies whether the generated code aligns with the problem description and then verifies whether the obtained result is consistent with the problem description. This step-by-step verification strategy significantly improves the verification process in self-correction. The excellent performance of our method on the challenging mathematical problem tasks, such as MATH and GSM8k, clearly demonstrates its effectiveness compared to the existing SFT-based methods (Yu et al., 2024; An et al., 2023).
@@ 3.1 Overview
We propose a Chain of Self-Correction (CoSC) to address mathematical questions in a self-correction manner. Given a mathematical question denoted as $q$, the CoSC mechanism initiates the process by generating a program $p$ that comprises a Python function. The program $p$ is then executed using a Python interpreter to produce an output $o$. However, the program $p$ may contain bugs or incorrect intermediate code due to potential misinterpretations of the question $q$, leading to erroneous outputs. Unlike existing methods that either base their reasoning on incorrect intermediate content or halt reasoning upon encountering an error, the CoSC mechanism introduces a self-correction rationale during the reasoning process. After the generation of program $p$ and its output $o$, the CoSC model analyzes them and generates a verification $v$. This verification includes suggestions for improvements to ensure the consistency of the program $p$ and its output $o$ with the mathematical question $q$, respectively. Following this, the CoSC mechanism draws a conclusion $c$ based on these suggestions, which is used to either refine the program $p$ or generate the final answer. This process is repeated until a conclusive answer is reached, which can be summarized as $answer=povcpovc\cdots povc$. By employing this reasoning method, we can annotate public mathematical datasets and use the annotated multi-round, self-correction data to fine-tune LLMs.
@@ 3.2.1 CoSC Foundational Learning
Existing mathematical reasoning datasets, such as chain-of-thoughts (CoT) (Wei et al., 2022) and program-of-thoughts (PoT) (Chen et al., 2023a), primarily contain single-round annotations without multi-round, self-correction solutions for the reasoning process. This makes it challenging to fine-tune models to inherently possess self-correction capabilities. To address this issue, we utilize GPT-4 ($\mathcal{G}$) and a few-shot approach to annotate questions from publicly available mathematical datasets, including MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021), to generate seeding data with our Chain of Self-Correction (CoSC) mechanism. Table 1 compares our seeding data of CoSC with recently proposed mathematical reasoning datasets. Detailed instructions and example cases of our CoSC can be found in Appendix A and Appendix B.
@@
Table 1: The comparison to the previous mathematical reasoning datasets. Note that our CoSC only employs questions from the original training set of MATH and GSM8K datasets. Particularly, we provide the number of annotation data as #Annotation in each mathematical reasoning method. Additionally, Correction signifies whether the self-correction is to be utilized.
Methods #Annotation Tool Correction LLM Used Source
RFT (Yuan et al., 2023) $>$100k LLaMA-2 GSM8K
Open-Platypus (Lee et al., 2023) 25k GPT-4 11 datasets with MATH
WizardMath (Luo et al., 2023) $>$96k ChatGPT MATH & GSM8K
Lila (Mishra et al., 2022) 134k - 20 datasets with MATH & GSM8K
MathInstruct (Yue et al., 2023) 260k GPT-4 14 datasets with MATH & GSM8K
ToRA (Gou et al., 2023b) 16k GPT-4 MATH & GSM8K
CoSC (Ours) 37k GPT-4 MATH & GSM8K
@@
With CoSC, we use GPT-4 to annotate the MATH and GSM8K training datasets. To explore diverse data, we apply nucleus sampling (Holtzman et al., 2019) during GPT-4 annotation. Each question is sampled in 3 times. However, for some complex questions, we can only get less than 1 correct solutions. For these problems, we apply 10 more samplings and retained up to 4 correct data. Finally, we filtered out incorrect answers and constructed 37k pieces of data using GPT-4. The algorithm of the generation with our CoSC for each question from the training set is shown in Algorithm 1.
@@
1:question $q$, model $\mathcal{G}$, prompt Weierstrass-$\wp$, external tools $\mathcal{E}$, stop condition Stop($\cdot$), iteration rounds $n_{q}$
2:$\tau_{0}\leftarrow\text{""}$" $\triangleright$ Trajectory Initialization
3:for $i\leftarrow 1$ to $n_{q}$ do
4:     $p_{i}\sim\mathbb{P}_{\mathcal{G}}(\cdot|\wp\oplus q\oplus\tau_{i-1})$triangleright Program Generation (Eq. 1)
5:     $o_{i}\leftarrow\mathcal{E}(p_{i})$triangleright Tool Execution
6:     $v_{i}\oplus c_{i}\sim\mathbb{P}_{\mathcal{G}}(\cdot|\wp\oplus q\oplus\tau_{i-1}\oplus p_{i}\oplus o_{i})$triangleright Verification and Conclusion Generation (Eq. 2)
7:     $\tau_{i}\leftarrow\tau_{i-1}\oplus p_{i}\oplus o_{i}\oplus v_{i}\oplus c_{i}$triangleright Trajectory Adjustment (Eq. 3)
8:     if Stop($c_{i}$then $\triangleright$ Stopping Criteria
9:         return $\tau_{i}$
10:     end if
11:end for
12:return $\tau_{n}$

### v2

@@ Abstract
Accurate mathematical reasoning with Large Language Models (LLMs) is crucial in revolutionizing domains that heavily rely on such reasoning. However, LLMs often encounter difficulties in certain aspects of mathematical reasoning, leading to flawed reasoning and erroneous results. To mitigate these issues, we introduce a novel mechanism, the Chain of Self-Correction (CoSC), specifically designed to embed self-correction as an inherent ability in LLMs, enabling them to validate and rectify their own results. The CoSC mechanism operates through a sequence of self-correction stages. In each stage, the LLMs generate a program to address a given problem, execute this program using program-based tools to obtain an output, subsequently verify this output. Based on the verification, the LLMs either proceed to the next correction stage or finalize the answer. This iterative self-correction process allows the LLMs to refine its reasoning steps and improve the accuracy of its mathematical reasoning. We implement CoSC using a two-phase fine-tuning approach. First, LLMs are trained with a relatively small volume of seeding data generated from GPT-4. Then, we enhance CoSC by training with a larger volume of self-generated data, without relying on GPT-4. Experiments show that CoSC significantly boosts performance on standard mathematical datasets compared to existing open-source LLMs. Notably, our CoSC-Code-34B model achieved a 53.5% score on the challenging MATH dataset, outperforming models like ChatGPT, GPT-4, and multi-modal LLMs such as GPT-4V and Gemini-1.0. Importantly, CoSC operates in a zero-shot manner without requiring demonstrations.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 (OpenAI, 2023a), have recently demonstrated state-of-the-art performance across a variety of natural language processing (NLP) tasks, including natural language generation and understanding (Chowdhery et al., 2023; Team et al., 2023; Anil et al., 2023; Penedo et al., 2023; Gao et al., 2024a, b). Despite their success, LLMs often struggle with mathematical reasoning tasks due to their lack of explicit logical reasoning and judgment, which are crucial for solving such problems. Moreover, there is a fundamental gap between natural language and the language of mathematical formulas, which further complicates these tasks. As a result, accurate mathematical reasoning remains an essential yet challenging capability for LLMs to develop, in order to further advance various domains. Consequently, it is still an open challenge to tackle mathematical problems for existing open-source LLMs (Touvron et al., 2023).
Figure 1: Comparison of four reasoning frameworks for solving an example mathematical question. (a) Chain of Thoughts (CoT) (Wei et al., 2022). (b) Program of Thoughts (PoT) (Chen et al., 2023a). (c) ToRA (Gou et al., 2023b) that incorporates CoT, PoT, and the utilization of tools. (d) Our proposed CoSC consists of a sequence of multiple self-correction stages (two stages are shown in this example). Each stage has four sub-stages: (p1) LLMs generate program w.r.t. the question; (o1) execute the program to obtain program output; (v1) perform two-step verification for consistency of the question with both the generated program and the program output; (c1) conclude a refined answer or continue the next subsequent self-correction stage depending on the verification result. The final answer is extracted from the last conclusion sub-stage with regular expression matching.
@@
Most current methods (Yu et al., 2023; Gou et al., 2023b) generate finetuning datasets by prompting GPT-4 to rephrase mathematical questions from various perspectives or incorporate chain-of-thoughts (CoT) analysis (Wei et al., 2022) and program-of-thoughts (PoT) code (Chen et al., 2023a) to diversify mathematical responses. As a result, LLMs trained on these datasets can comprehend different questions and learn to use code to solve mathematical problems. However, precise multi-round reasoning capabilities remain challenging for them. Once potential errors occur throughout reasoning stages, it can lead to incorrect results. Consequently, it is essential to incorporate a self-correction mechanism into mathematical responses, which can enable LLMs to learn to correct themselves in multiple rounds. Besides, while supervised finetuning datasets can be developed with GPT-4’s assistance, it still requires human experiments with different prompts and the cost of using the interface. This highlights the need for effective finetuning with unlabeled datasets.
@@
Using GPT-4 can be expensive, especially when dealing with large volumes of training data. Alternatively, we propose a cost-free method to further boost performance through self enhancement in the second phase. The second phase, referred to as CoSC self enhancement, builds upon foundational learning by further adapting the LLMs obtained from the first phase with self-generated trajectories. These trajectories are produced by the models trained in the foundational phase, allowing for the generation of a substantial volume of data without GPT-4 intervention. In both phases, we retain only the trajectories whose answers match the ground-truth labels of the corresponding questions.
@@
  • To implement the CoSC mechanism at a low cost, we introduce a two-phase finetuning approach. The first phase involves CoSC foundational learning, where we use a relatively small volume of seeding data generated by GPT-4. In the second phase, CoSC self enhancement occurs using a larger volume of self-generated data with the model obtained from the first phase, without relying on paid GPT-4.
  • Our comprehensive experiments demonstrate that the CoSC mechanism provides a new benchmark for performance on established mathematical datasets when compared to existing open-source LLMs. Notably, our CoSC-Code-34B model achieves superior performance over both closed-source LLMs and some of multi-modal LLMs, particularly on the challenging MATH dataset.
  • The proposed CoSC mechanism, by embedding self-correction as an inherent capability in LLMs, enables them to think before responding to a question, creating an internal chain of self-correction to progressively verify and rectify their original answers. It is more akin to the slow thinking process of humans, which is particularly helpful in solving difficult mathematical reasoning problems. This approach can provide valuable insights for future research and contribute to the ongoing advancement of LLMs.
@@ 2.1 LLMs for Mathematical Reasoning
Mathematical reasoning (Liu et al., 2023a; Wang et al., 2023a; Huang et al., 2024b; Toshniwal et al., 2024; Chen et al., 2024a; Zhang et al., 2024) is a challenging reasoning task for LLMs, which requires the ability to understand mathematical concepts, computation and multi-round reasoning. Existing mathematical reasoning approaches can be broadly classified into three categories: (1) Prompting methods (Wei et al., 2022; Chen et al., 2023a; Wang et al., 2023b) focus on extracting the inherent mathematical reasoning skills of LLMs by utilizing well-crafted prompting strategies during inference. Notably, they leverage the existing knowledge in LLMs without the need for parameter updates. (2) Pretraining methods (Azerbayev et al., 2023; Fu et al., 2023; Yang et al., 2024) pre-train LLMs on large-scale corpora containing mathematical problems and related content with language modeling objectives. The goal is to train a base foundation language model as a platform for mathematical domain. (3) Finetuning methods (Luo et al., 2023; Yu et al., 2023; Yue et al., 2023; Gou et al., 2023b; Liu et al., 2023b; Gao et al., 2024c; Mitra et al., 2024) refine the mathematical reasoning capabilities of LLMs by offering more targeted training. Central to this approach is the generation of high-quality question-response pair datasets, which needs the assistance of complexity-based chain-of-thoughts prompting or tools-based augmentation. In this paper, we propose a Chain of Self-Correction (CoSC) along this direction that incorporates an iterative self-correction process into datasets generation.
@@ 2.2 Existing methods related to Self-Correction in LLMs
Mathematical reasoning poses a significant challenge due to its demand for precise multi-round logical reasoning to solve problems. The potential for errors increases with each reasoning step, making it crucial for LLMs to have the ability to self-correct in order to produce accurate results. There are some recent studies (Chen et al., 2023b; Gou et al., 2023a; Lightman et al., 2023; Huang et al., 2024a; Chen et al., 2024b) attempt to enable large language models to perform self-correction by either prompting methods or fine-tuning methods. For prompting methods, they can correct their responses when interacting with external tools (Gou et al., 2023a; Chen et al., 2023b), such as search engines and calculators, or designing complex prompts (Chen et al., 2024b). Notably, prompting methods can be orthogonally combined with finetuning methods. For finetuning methods, previous works (Yu et al., 2024; An et al., 2023) only model single-round correction during the training stage, and perform verification in a straightforward manner during the inference stage.
@@ 2.3 Difference between the proposed method and existing self-correction techniques
The existing self-correction research can be broadly classified into two categories: prompt-based methods (Gou et al., 2023a; Chen et al., 2023b, 2024b) and SFT-based methods (Yu et al., 2024; An et al., 2023). Our approach falls under the SFT-based category and fundamentally differs from prompt-based methods. Specifically, prompt-based methods leverage the intrinsic capabilities of LLMs for mathematical reasoning. In contrast, our method embeds the chain of self-correction, a strong reasoning mechanism, as an inherent capability in the LLMs through parameter fine-tuning. Furthermore, compared to existing SFT-based methods, our approach models mathematical reasoning as a multi-round procedure during the training stage. In addition, these SFT-based works (Yu et al., 2024; An et al., 2023) perform the verification process in a straightforward manner, whereas our work conducts verification in a step-by-step manner by generating intermediate verification steps. Similar to the essence of the CoT approach, our model first verifies whether the generated code aligns with the problem description and then verifies whether the obtained result is consistent with the problem description. This step-by-step verification strategy significantly improves the verification process in self-correction. The excellent performance of our method on the challenging mathematical problem tasks, such as MATH and GSM8k, clearly demonstrates its effectiveness compared to the existing SFT-based methods (Yu et al., 2024; An et al., 2023).
@@ 3.1 Overview
We propose a Chain of Self-Correction (CoSC) to address mathematical questions in a self-correction manner. Given a mathematical question denoted as $q$, the CoSC mechanism initiates the process by generating a program $p$ that comprises a Python function. The program $p$ is then executed using a Python interpreter to produce an output $o$. However, the program $p$ may contain bugs or incorrect intermediate code due to potential misinterpretations of the question $q$, leading to erroneous outputs. Unlike existing methods that either base their reasoning on incorrect intermediate content or halt reasoning upon encountering an error, the CoSC mechanism introduces a self-correction rationale during the reasoning process. After the generation of program $p$ and its output $o$, the CoSC model analyzes them and generates a verification $v$. This verification includes suggestions for improvements to ensure the consistency of the program $p$ and its output $o$ with the mathematical question $q$, respectively. Following this, the CoSC mechanism draws a conclusion $c$ based on these suggestions, which is used to either refine the program $p$ or generate the final answer. This process is repeated until a conclusive answer or a maximum limit of self-correction stages is reached, which can be summarized as $answer=povcpovc\cdots povc$. By employing this reasoning method, we can annotate public mathematical datasets and use the annotated multi-round, self-correction data to fine-tune LLMs.
@@ 3.2.1 CoSC Foundational Learning
Existing mathematical reasoning datasets, such as chain-of-thoughts (CoT) (Wei et al., 2022) and program-of-thoughts (PoT) (Chen et al., 2023a), primarily contain single-round annotations without multi-round, self-correction solutions for the reasoning process. This makes it challenging to fine-tune models to inherently possess self-correction capabilities. To address this issue, we utilize GPT-4 ($\mathcal{G}$) and a few-shot approach to annotate questions from publicly available mathematical datasets, including MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021), to generate seeding data with our Chain of Self-Correction (CoSC) mechanism. Detailed instructions and example cases of our CoSC can be found in Appendix A and Appendix B.
@@
With CoSC, we use GPT-4 to annotate the MATH and GSM8K training datasets. To explore diverse data, we apply nucleus sampling (Holtzman et al., 2019) during GPT-4 annotation. Each question is sampled in 3 times. However, for some complex questions, we are unable to obtain even a single correct solution. For these problems, we apply 10 more samplings and retained up to 4 correct data. Finally, we filtered out incorrect answers and constructed 37k pieces of data using GPT-4. The algorithm of the generation with our CoSC for each question from the training set is shown in Algorithm 1.
@@
Input: question $q$, model $\mathcal{G}$, prompt Weierstrass-$\wp$, external tools $\mathcal{E}$, stop condition Stop($\cdot$), iteration rounds $n_{q}$
1:  $\tau_{0}\leftarrow\texttt{""}$"
2:  for $i\leftarrow 1$ to $n_{q}$ do
3:    $p_{i}\sim\mathbb{P}_{\mathcal{G}}(\cdot|\wp\oplus q\oplus\tau_{i-1})$
4:    $o_{i}\leftarrow\mathcal{E}(p_{i})$
5:    $v_{i}\oplus c_{i}\sim\mathbb{P}_{\mathcal{G}}(\cdot|\wp\oplus q\oplus\tau_{i-1}\oplus p_{i}\oplus o_{i})$
6:    $\tau_{i}\leftarrow\tau_{i-1}\oplus p_{i}\oplus o_{i}\oplus v_{i}\oplus c_{i}$
7:    if Stop then
8:       return $\tau_{i}$
9:    end if
10:  end for
11:  return $\tau_{n}$